OpenStudy (anonymous):

Can anyone isolate y? 2y(x^2)+7x(y^2)=408

OpenStudy (anonymous):

yes have u tried distributing?

OpenStudy (anonymous):

i tried but i cant seem to get y to stay on one side

OpenStudy (anonymous):

i got this\[2xy+7y^2=\frac{ 408 }{ x }\]

OpenStudy (callisto):

2y(x^2)+7x(y^2)=408 7x(y^2)+2(x^2)y+7x(y^2)-408=0 Quadratic formula \[y=\frac{-(2x^2)\pm\sqrt{(2x^2)^2-4(7x)(-408)}}{2(7x)}\]

OpenStudy (anonymous):

oh crap i completely forgot about the quadratic formula dang it thank you

OpenStudy (callisto):

Second line should be 7x(y^2)+2(x^2)y+-408=0

OpenStudy (anonymous):

ok thank you

OpenStudy (callisto):

Welcome :)

OpenStudy (callisto):

Not sure if it is the right approach though.

OpenStudy (anonymous):

Firstly, there is no way of "isolating" y here. You can solve for 'y' given some value of 'x' using the quadratic formula like @Callisto suggested. @Nogarenshi

OpenStudy (anonymous):

well its the only one i got so i will go for it this is for multivariable calculus so i forgot about the quadratic formula

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

um i am looking at this again and something doesnt seem right can anyone else clarify.

OpenStudy (anonymous):

What doesn't seem right?

OpenStudy (anonymous):

first open up the bracket: 2y(x^2)+7x(y^2)=408 by opening up you'll have something similar to this 2yx^2+7xy^2=408 then factor out y:. y(2x^2+7x^2)=408 finally divide both side by (2x^2+7x^2) and your answer will be Y=408/2x^2+7x^2

OpenStudy (anonymous):

@hassannoor0480 That's not how it works. By factoring out 'y', there is still a 'y' left with the "7xy". Like I mentioned, there is no way of isolating y here.

OpenStudy (anonymous):

ok actually nevermind what castillo said works i notice i wrote something down wrong

OpenStudy (anonymous):

Yeah, what Callisto says is true, he isn't factoring anything out:\[ ~~~~~\color{red}{(7x)}\cdot y^2+~\color{green}{(2x^2)}\cdot y ~~\color{blue}{- 408} = 0 \\ a = \color{red}{7x},~~~~~~b=\color{green}{2x^2},~c=\color{blue}{-408}\\ \]

OpenStudy (anonymous):

yeah genius am wrong i thought y is not squared